# Jianxiong Gao, Yichang Liu, Baofeng Yang, Jianfeng Feng, Yanwei Fu

### v1

@@ Abstract
In this paper, we introduce CineBrain, the first large-scale dataset featuring simultaneous EEG and fMRI recordings during dynamic audiovisual stimulation. Recognizing the complementary strengths of EEG’s high temporal resolution and fMRI’s deep-brain spatial coverage, CineBrain provides approximately six hours of narrative-driven content from the popular television series The Big Bang Theory for each of six participants. Building upon this unique dataset, we propose CineSync, an innovative multimodal decoding framework integrates a Multi-Modal Fusion Encoder with a diffusion-based Neural Latent Decoder. Our approach effectively fuses EEG and fMRI signals, significantly improving the reconstruction quality of complex audiovisual stimuli. To facilitate rigorous evaluation, we introduce Cine-Benchmark, a comprehensive evaluation protocol that assesses reconstructions across semantic and perceptual dimensions. Experimental results demonstrate that CineSync achieves state-of-the-art video reconstruction performance and highlight our initial success in combining fMRI and EEG for reconstructing both video and audio stimuli. Project Page: https://jianxgao.github.io/CineBrain.
@@
footnotetext: Email: jxgao22@m.fudan.edu.cn, yanweifu@fudan.edu.cn.
@@ 1 Introduction
The human brain, often described as the most complex and enigmatic organ, has long intrigued researchers. Functional Magnetic Resonance Imaging (fMRI) and Electroencephalography (EEG) are two widely employed non-invasive techniques to capture neural responses, facilitating the decoding of brain activities evoked by various stimuli. While fMRI measures blood oxygenation changes in the brain induced by neuronal activity, EEG captures electrical signals resulting directly from neuronal firing. Previous research [22, 47, 6, 13, 14] has demonstrated that both modalities effectively decode stimuli encompassing visual, linguistic, and auditory information.
However, these modalities offer complementary strengths while also suffering from intrinsic limitations. Specifically, fMRI can probe deep-brain neural activities but is hindered by relatively low temporal resolution, potentially overlooking transient but critical neural dynamics. Conversely, EEG provides superior temporal resolution suitable for capturing rapid neural oscillations, yet its signals predominantly reflect superficial cortical activities. Consequently, decoding continuous neural responses, such as those elicited by audiovisual stimuli, solely through either modality inevitably results in significant informational loss. Motivated by these complementary characteristics, a natural question emerges: Can we leverage the high temporal resolution of EEG to compensate for the inherent temporal limitations of fMRI? This paper presents an initial exploration into addressing this challenge by reconstructing video and audio from human brain activity.
Currently, no publicly available dataset provides simultaneous EEG and fMRI recordings during temporally dynamic stimulation such as video or audio presentations. To bridge this gap, we introduce the cureated CineBrain, the first large-scale multimodal stimulation dataset that captures synchronized fMRI and EEG signals.
Literature [18, 16, 43] highlights that engaging narratives can naturally capture audience attention, sustain neural engagement, and provide deep insights into complex brain dynamics. Additionally, neuroscientific findings [33, 41] emphasize that different regions of the human brain are highly interconnected and interact dynamically during perceptual processing. Motivated by these insights, we designed the CineBrain dataset to incorporate audiovisual stimuli closely resembling real-world experiences, where participants simultaneously watched and listened to dynamic content. Specifically, we selected episodes from the popular human-centric television series The Big Bang Theory, providing approximately six hours of narrative-driven audiovisual content for each of six participants. Recordings were conducted in a 3T MRI scanner, with participants wearing customized non-magnetic headphones and EEG caps. As illustrated in Fig. 2, participants 1–4 exhibited consistent activation patterns in both fMRI and EEG modalities when presented with identical stimuli. Further details about the CineBrain dataset are provided in Sec. 3.
Leveraging the CineBrain dataset, we propose an innovative framework called CineSync, aiming to reconstruct both video and audio stimuli from brain signals. We introduce a Multi-Modal Fusion Encoder that effectively integrates information from fMRI and EEG. Initially, the encoder applies a dual transformer-based architecture to separately encode multi-frame signals from each modality. To ensure accurate and meaningful multimodal feature integration, both modalities are jointly aligned with visual and textual representations using a combined contrastive loss. Subsequently, these fused brain representations serve as inputs to our Neural Latent Decoder. To guarantee both semantic coherence and perceptual fidelity in reconstructions, we train a diffusion-based Neural Latent Decoder that integrates brain-derived features with latent noise.
Given that CineBrain stimuli are human-centric audiovisual materials, reconstructing such complex content is challenging even for advanced generative models. Therefore, selecting robust and suitable evaluation metrics is crucial. To address this need, we introduce the Cine-Benchmark, which comprehensively assesses video reconstructions on semantic and perceptual dimensions, as well as audio reconstructions. Experimental results confirm that our approach achieves state-of-the-art performance in brain-signal-based visual reconstruction, particularly by effectively integrating EEG and fMRI modalities to improve the reconstruction quality of temporally dynamic stimuli within the challenging CineBrain dataset.
@@
  • For the first time, we provide an initial exploration and validation demonstrating that combining the complementary strengths of fMRI and EEG significantly enhances the reconstruction quality of temporally dynamic stimuli.
  • We introduce CineBrain, the first multimodal audiovisual stimulation dataset featuring meticulously paired modalities, including fMRI, EEG, video, audio, and captions. CineBrain provides challenging tasks suitable for diverse experimental setups and fosters research in multimodal brain decoding.
  • We propose CineSync, a versatile multimodal brain decoding framework comprising a Multi-Modal Fusion Encoder and a diffusion-based Neural Latent Decoder. CineSync achieves state-of-the-art video reconstruction performance and can be readily adapted for audio reconstruction tasks.
  • We develop Cine-Benchmark, a comprehensive evaluation protocol designed to effectively assess the reconstruction quality of video and audio from brain signals.
@@ 2.1 Neuro-Stimulus Dataset
Data is fundamental to deep learning. To facilitate neural decoding experiments, various datasets have been proposed. For visual stimuli, NSD dataset [1] is a widely used large-scale fMRI dataset collected from image-based stimuli, while Wen dataset [44] and fMRI-Video-WebVid [27] contain fMRI data acquired from video stimuli. Additionally, the fMRI-Shape [13] and fMRI-Objaverse [14] datasets provide fMRI recordings from 3D video stimuli. For EEG data, the SEED-DV [30] dataset includes EEG recordings from video stimuli. Tang’s dataset [42] focuses specifically on auditory stimulation. However, these existing datasets are limited because each dataset only captures single-modality brain signals responding to single-modality stimuli. To address this limitation, in this paper, we propose a multimodal audiovisual stimulus dataset named CineBrain and conduct experiments based on it.
Wen [44] Video 3 0/3 3.07 h 5520
fMRI-Video-FCVID [27] Video 3 1/1 1.11 h 400
fMRI-Video-WebVid [27] Video 5 2/2 2.67 h 1200
SEED-DV [30] Video 20 10/10 0.76 h 1400
CineBrain (Ours) Audio+Video 6 2/4 6 h 5400
@@ 2.2 Neural Decoding
Neural decoding refers to the process of reconstructing external stimuli perceived by the human brain from brain signals that reflect underlying neural activity. This process poses significant challenges for both encoding brain signals and decoding the corresponding stimuli.
Functional magnetic resonance imaging (fMRI), known for its excellent spatial resolution and non-invasive characteristics, has demonstrated effectiveness in reconstructing visual stimuli. Methods such as MinD-Vis [6] and Mind-Eye [39] have successfully achieved image reconstruction. Techniques including Mind-Video [7], GLFA [27], and NeuroClips [15] have further extended these accomplishments into video reconstruction. More recently, MinD-3D [13] and MinD-3D++ [14] have explored reconstructing three-dimensional visual content. Electroencephalography (EEG), characterized by high temporal resolution and continuous signals, has also shown promise in neural decoding. EEG-based methods have successfully reconstructed various stimuli, including videos [31], 3D structures [17], and auditory signals [23, 8].
However, existing methods have not considered combining different types of brain signals, such as fMRI and EEG, which possess significantly different characteristics. In this paper, we aim to explore the effectiveness of integrating fMRI and EEG signals to enhance stimulus reconstruction.
@@ 2.3 Diffusion Models
Diffusion models [20, 38, 35] are a powerful generative framework known for their capability to produce high-quality images and can also be extended to other data modalities. The fundamental principle involves defining a forward diffusion process, in which clean data is progressively corrupted by adding Gaussian noise. A neural network, typically a U-Net architecture, is then trained to reverse this diffusion process, iteratively removing the noise to reconstruct high-quality data. DiT [34] has successfully scaled up diffusion models through transformer-based architectures, significantly improving their generative capabilities. Consequently, DiT has been effectively applied across multiple modalities, including text-to-image generation [34, 10, 25], text-to-video generation [45, 19, 3], and text-to-audio generation [5, 9]. In our study, motivated by the strong distribution modeling capabilities of diffusion models, we leverage diffusion models specifically trained on video and audio data to reconstruct these modalities from multimodal brain signals, including fMRI and EEG.
@@ 3.1 Subjects and Experiments
We select episodes from the popular TV series The Big Bang Theory as audiovisual stimuli due to its engaging nature, relevance to everyday experiences, and its effectiveness in sustaining participants’ attention over the 18-minute viewing periods. To enhance the diversity of stimuli, participants collectively view 20 different episodes. All participants watch the first 10 episodes of Season 7; additionally, Participants 1, 2, and 6 watch 10 episodes from Season 9, while the remaining participants view 10 episodes from Season 11.
During the experiment, the video resolution is downsampled to 720p to match the MRI’s in-bore LED screen. Episodes of varying lengths are standardized to 18 minutes of viewing time, easing subsequent data processing. Each 18-minute viewing constitutes one “run,” yielding 20 runs per participant. To maintain data quality and reduce fatigue, runs are grouped into sessions of two or three episodes each, with no more than five runs conducted on a single day. In total, each participant contributes approximately 6 hours of concurrent fMRI and EEG alongside ECG.
@@ 3.2 Data Acquisition and Preprocessing
fMRI. The fMRI acquisition protocol follows established experimental standards [13, 14, 27]. A 3T scanner equipped with a 32-channel RF head coil is employed to obtain high-resolution T1-weighted structural images and functional data for the decoding experiments.
T1-weighted images were obtained using an MPRAGE sequence (0.8-mm isotropic resolution, TR = 2500 ms, TE = 2.22 ms, flip angle $8^{\circ}$). Functional data were recorded using a gradient-echo EPI sequence with whole-brain coverage (2-mm isotropic resolution, TR = 800 ms, TE = 37 ms, flip angle $52^{\circ}$, multi-band factor = 8). With a sampling frequency of 1.25 Hz, each video run generated 1350 functional MRI frames.
Preprocessing is conducted using the widely adopted fMRIPrep pipeline [12, 11]. Unlike prior studies focused solely on visual decoding [13, 6, 7], our experimental design also incorporated auditory stimuli, leading us to select both visual and auditory regions of interest (ROIs).
@@
For the selection of visual regions of interest (ROIs), we follow the regions commonly reported in [13, 14]. The ROIs were defined using the parcellation provided by the Human Connectome Project Multi-Modal Parcellation (HCP-MMP) in the 32k_fs_LR space. These ROIs included areas such as “V1, V2, V3, V3A, V3B, V3CD, V4, LO1, LO2, LO3, PIT, V4t, V6, V6A, V7, V8, PH, FFC, IP0, MT, MST, FST, VVC, VMV1, VMV2, VMV3, PHA1, PHA2, PHA3”. For the auditory cortex, we selected regions corresponding to “4, 7AL, 7Am, 7m, 7PC, 7PL, 7Pm, 8Ad, 8Av, 8BM, 8C, 9a, 9p, 10d, 10v, 44, 45, 46, 47m, A1, IPS1, p32, PGp, s32, STGa, STSda, STSdp, STSva, STSvp, TPOJ1, TPOJ2, TPOJ3”. The selection of these ROIs is illustrated in Fig. 3. The visual ROIs contain 8,405 voxels, and the auditory ROIs contain 10,541 voxels.
EEG. EEG data are captured using an MRI-compatible 64-channel cap at 1000 Hz, simultaneously recording ECG signals. Precise synchronization between EEG and fMRI data is ensured by logging fMRI TR timings. EEG preprocessing involves a multi-step artifact removal approach, targeting scanner-induced noise and biological interference while preserving neural activity. This includes bandpass filtering (0.1–30 Hz) to remove baseline drift and muscle artifacts, and a 50 Hz notch filter to reduce powerline interference [4, 26, 28, 29]. ECG artifacts are initially mitigated using QRS-based techniques, followed by independent component analysis (ICA) to isolate residual artifacts. ECG recordings enable correlation analyses to adaptively refine artifact removal, culminating in clean EEG data suitable for subsequent analyses.
Video Stimuli. Original video stimuli consist of twenty 18-minute episodes at 24 fps and 1080p. For our experiments, we resize and crop these videos to a resolution of 480$\times$720 and segment each into 4-second clips (33 frames per clip). Thus, each participant generates 5400 clips, partitioned into 4860 training clips (first 18 episodes) and 540 testing clips (final 2 episodes).
Audio. Audio from each video episode is segmented into corresponding 4-second clips, resulting in the same number of samples as the video clips, maintaining consistency for training and evaluation.
Text Descriptions. To support multimodal decoding experiments, textual descriptions are generated separately for each audio and video clip. Video descriptions are generated using vision-language models (Qwen2.5-VL [2] and Llava-Video [46]) to facilitate contrastive learning. Audio clips are transcribed using Whisper-large-v3 [37], providing rich textual inputs aligned with auditory data for enhanced multimodal analyses.

### v2

# Jianxiong Gao1, Yichang Liu1, Baofeng Yang1, Jianfeng Feng1, Yanwei Fu1,2

@@ Abstract
Most research decoding brain signals into images, often using them as priors for generative models, has focused only on visual content. This overlooks the brain’s natural ability to integrate auditory and visual information, for instance, sound strongly influences how we perceive visual scenes. To investigate this, we propose a new task of reconstructing continuous video stimuli from multimodal brain signals recorded during audiovisual stimulation. To enable this, we introduce CineBrain, the first large-scale dataset that synchronizes fMRI and EEG during audiovisual viewing, featuring six hours of The Big Bang Theory episodes for cross-modal alignment. We also conduct the first systematic exploration of combining fMRI and EEG for video reconstruction and present CineSync, a framework for reconstructing dynamic video using a Multi-Modal Fusion Encoder and a Neural Latent Decoder. CineSync achieves state-of-the-art performance in dynamic reconstruction, leveraging the complementary strengths of fMRI and EEG to improve visual fidelity. Our analysis shows that auditory cortical activations enhance decoding accuracy, highlighting the role of auditory input in visual perception. Project Page: https://jianxgao.github.io/CineBrain.
@@ 1 Introduction
In the vision community, considerable effort has been devoted to decoding brain signals—particularly fMRI—into images [chen2023seeing, mind-eye, scotti2024mindeye2sharedsubjectmodelsenable], videos [mind-video, lichong_video], and even 3D shapes [mind3d], often using neural signals as priors to guide visual generative models. While this prior-guided approach is widely adopted, it largely reduces neural decoding to a tool for controlling generative models rather than probing the principles of human perception. In contrast, the human brain, arguably the most intricate biological system, excels at integrating information across multiple sensory modalities. Humans naturally perceive the world through rich audiovisual experiences, seamlessly combining visual and auditory cues to form coherent, dynamic representations of their surroundings. A classic example is the McGurk effect [mcgurk_nature], where conflicting auditory and visual speech cues produce a third, illusory percept, highlighting the fundamental interplay between vision and hearing. Beyond generative priors, understanding how the brain achieves such integration is central to bridging human cognition and machine perception.
Technically, most existing studies [chen2023seeing, mind-video, mind3d, eeg2video, eeg3d] in neural decoding focus on reconstructing visual content from visual-only stimuli, typically relying on a single neural modality such as fMRI or EEG. This limited scope overlooks two critical aspects of perception. First, auditory signals strongly modulate visual processing [sporns2007brain, park2013structural], shaping attention, emotion, and context during naturalistic viewing. Second, fMRI and EEG provide complementary insights: fMRI offers fine-grained spatial localization, while EEG captures millisecond-level temporal dynamics. Ignoring the cross-modal and cross-modality interactions restricts the fidelity and realism of reconstructed visual stimuli.
To address this gap, we propose both a new task and a novel large-scale dataset. Specifically, we present the task of reconstructing continuous video stimuli from multimodal brain signals recorded during naturalistic audiovisual stimulation. This task requires capturing both the spatial and temporal aspects of neural encoding, enabling exploration of how auditory input influences visual perception. However, progress in this area has been limited by the lack of datasets that synchronize fMRI and EEG recordings under ecologically valid, temporally dynamic audiovisual conditions. To overcome this challenge, we present CineBrain, the first large-scale multimodal dataset that synchronizes fMRI and EEG during naturalistic audiovisual viewing. Drawing inspiration from neurocinematics research [hasson2008neurocinematics, watabe2017neural], which demonstrates that narrative-driven content naturally sustains attention and elicits complex brain dynamics, CineBrain includes approximately six hours of The Big Bang Theory episodes, viewed by six participants under 3T fMRI scanning with concurrent EEG recordings captured using custom non-magnetic equipment. As illustrated in Fig. 2, the participants exhibit highly consistent activation patterns across both modalities, providing strong cross-modal alignment that is critical for multimodal decoding tasks.
Building on CineBrain, we conduct the first systematic exploration of fMRI–EEG integration for video reconstruction and present CineSync, a novel framework for reconstructing dynamic visual stimuli from multimodal brain signals. CineSync employs a Multi-Modal Fusion Encoder with a dual-transformer design to independently process fMRI and EEG sequences before merging them via a learned fusion projector. To achieve semantic alignment across modalities, brain representations are anchored to visual and textual embeddings using a joint contrastive objective. The resulting multimodal embeddings are decoded by a Neural Latent Decoder, a diffusion-based model that fuses brain-derived features with latent noise to generate semantically coherent and perceptually realistic videos.
Evaluation using semantic and perceptual metrics demonstrates that CineSync achieves state-of-the-art performance on temporally dynamic reconstruction within CineBrain. It effectively leverages the complementary strengths of fMRI and EEG to enhance visual fidelity. Notably, auditory-related cortical activations improve decoding accuracy, highlighting cross-modal facilitation in natural perception, and ablation studies show that increasing EEG representational capacity further boosts reconstruction quality, underscoring EEG’s critical role in capturing rapid neural dynamics for fine-grained video reconstruction.
@@
  • A novel task: We introduce the challenge of reconstructing continuous video stimuli from multimodal brain signals, modeling spatial and temporal neural encoding to reveal how auditory input influences visual perception.
  • A new dataset: We present CineBrain, the first large-scale multimodal dataset that synchronizes fMRI and EEG during naturalistic audiovisual viewing, providing critical cross-modal alignment for decoding tasks.
  • A new framework: We propose CineSync, a framework for reconstructing dynamic video stimuli using a dual-transformer-based Multi-Modal Fusion Encoder (MFE) and a diffusion-based Neural Latent Decoder.
  • State-of-the-art performance: CineSync achieves superior reconstruction quality, leveraging fMRI–EEG complementarity, and highlighting the role of auditory-related cortical activations in improving decoding accuracy.
@@ 2 Related Work
Wen [wen_video] Video 3 0/3 3.07 h 5520
fMRI-Video-FCVID [lichong_video] Video 3 1/1 1.11 h 400
fMRI-Video-WebVid [lichong_video] Video 5 2/2 2.67 h 1200
SEED-DV [SEED-DV] Video 20 10/10 0.76 h 1400
CineBrain (Ours) Audio+Video 6 2/4 6 h 5400
@@ 2.1 Neuro-Stimulus Dataset
Data is fundamental to deep learning. To facilitate neural decoding experiments, various datasets have been proposed. For visual stimuli, NSD dataset [NSD] is a widely used large-scale fMRI dataset collected from image-based stimuli, while Wen dataset [wen_video] and fMRI-Video-WebVid [lichong_video] contain fMRI data acquired from video stimuli. Additionally, the fMRI-Shape [mind3d] and fMRI-Objaverse [mind3d++] datasets provide fMRI recordings from 3D video stimuli. For EEG data, the SEED-DV [SEED-DV] dataset includes EEG recordings from video stimuli. Tang’s dataset [tang_audio] focuses specifically on auditory stimulation. However, these existing datasets are limited because each dataset only captures single-modality brain signals responding to single-modality stimuli. To address this limitation, in this paper, we propose a multimodal audiovisual stimulus dataset named CineBrain and conduct experiments based on it.
@@ 2.2 Neural Decoding
Neural decoding aims to reconstruct external stimuli perceived by the human brain from recorded neural signals. This task remains challenging due to the complex nature of both neural encoding and stimulus reconstruction. Functional magnetic resonance imaging (fMRI), with its high spatial resolution and non-invasive acquisition, has shown strong performance in visual reconstruction. Recent works have achieved notable results in image reconstruction [chen2023seeing, mind-eye, sun2023contrastattenddiffusedecode] and extended these advances to video [mind-video, lichong_video, sun2025neuralflix, gong2024neuroclips, wang2025neuronsemulatinghumanvisual] and 3D reconstruction [mind3d, mind3d++]. Electroencephalography (EEG), offering millisecond-level temporal resolution and continuous recording, has also proven effective for neural decoding. EEG-based methods have reconstructed images [singh2023eeg2imageimagereconstructioneeg, bai2023dreamdiffusiongeneratinghighqualityimages, song2024decoding], videos [eeg2video, liu2025dynamindreconstructingdynamicvisual], 3D structures [eeg3d]. Despite these advances, most existing works have focused on single-modality decoding. The complementary characteristics of fMRI and EEG, where fMRI provides fine-grained spatial information and EEG offers precise temporal dynamics, have not yet been fully explored. In this paper, we investigate the integration of fMRI and EEG signals to exploit their complementary advantages and enhance the reconstruction of perceived stimuli.
@@ 2.3 Diffusion Models
Diffusion models [ddpm, ldm, sdxl] are a powerful generative framework known for their capability to produce high-quality images and can also be extended to other data modalities. The fundamental principle involves defining a forward diffusion process, in which clean data is progressively corrupted by adding Gaussian noise. A neural network, typically a U-Net architecture, is then trained to reverse this diffusion process, iteratively removing the noise to reconstruct high-quality data. DiT [DiT] has successfully scaled up diffusion models through transformer-based architectures, significantly improving their generative capabilities. Consequently, DiT has been effectively applied across multiple modalities, including text-to-image generation [DiT, sd3, flux2024], text-to-video generation [cogvideox, he2022lvdm, chen2024videocrafter2], and text-to-audio generation [f5tts, E2TTS]. In our study, motivated by the strong distribution modeling capabilities of diffusion models, we leverage diffusion models specifically trained on video and audio data to reconstruct these modalities from multimodal brain signals, including fMRI and EEG.
@@ 3.1 Subjects and Experiments
Since prolonged MRI scanning can cause discomfort and reduced attention, participants cannot stay in the scanner for long periods. Therefore, we select episodes from the popular TV series The Big Bang Theory as audiovisual stimuli. The show’s engaging narratives, familiarity with daily life, and rich audiovisual content help maintain participants’ attention throughout the 18-minute viewing sessions. To ensure diverse yet comparable stimuli, participants collectively view 20 different episodes. All participants watch the first 10 episodes of Season 7; in addition, Participants 1, 2, and 6 view 10 episodes from Season 9, while the remaining participants watch 10 episodes from Season 11.
During the experiment, the video resolution is downsampled to 720p to match the MRI’s in-bore LCD screen. Episodes of varying lengths are standardized to 18 minutes of viewing time, easing subsequent data processing. Each 18-minute viewing constitutes one “run”, yielding 20 runs per participant. To maintain data quality and reduce fatigue, runs are grouped into sessions of two or three episodes each, with no more than five runs conducted on a single day. In total, each participant contributes approximately 6 hours of concurrent fMRI and EEG alongside ECG.
@@ 3.2 Data Acquisition and Preprocessing
fMRI. The fMRI acquisition protocol follows established experimental standards [mind3d, mind3d++, lichong_video]. A 3T scanner equipped with a 32-channel RF head coil is employed to obtain high-resolution T1-weighted structural images and functional data for the decoding experiments. T1-weighted images were obtained using an MPRAGE sequence (0.8-mm isotropic resolution, TR = 2500 ms, TE = 2.22 ms, flip angle $8^{\circ}$). Functional data were recorded using a gradient-echo EPI sequence with whole-brain coverage (2-mm isotropic resolution, TR = 800 ms, TE = 37 ms, flip angle $52^{\circ}$, multi-band factor = 8). With a sampling frequency of 1.25 Hz, each video run generated 1350 functional MRI frames.
Preprocessing is performed using the widely adopted fMRIPrep pipeline [fmriprep1, fmriprep2]. Unlike previous studies that focused exclusively on visual decoding [mind3d, chen2023seeing, mind-video], our experimental design additionally incorporates auditory stimuli, prompting the selection of both visual and auditory regions of interest (ROIs). The selected ROIs are illustrated in Fig. 3. Specifically, the visual ROIs comprise 8,405 voxels, while the auditory ROIs comprise 10,541 voxels. Detailed ROI definitions are provided in the supplementary material.
@@
Figure 4: Architectural exploration for integrating fMRI and EEG. We compare five encoder variants, each adopting a distinct fusion mechanism to integrate the complementary information from fMRI and EEG signals.
EEG. EEG data are captured using an MRI-compatible 64-channel cap at 1000 Hz, simultaneously recording ECG signals. Precise synchronization between EEG and fMRI data is ensured by logging fMRI TR timings. EEG preprocessing involves a multi-step artifact removal approach, targeting scanner-induced noise and biological interference while preserving neural activity. This includes bandpass filtering (0.1–30 Hz) to remove baseline drift and muscle artifacts, and a 50 Hz notch filter to reduce powerline interference [chen2022dataset, lee2024eav, li2024visual, li2025neurobolt]. ECG artifacts are initially mitigated using QRS-based techniques, followed by independent component analysis to isolate residual artifacts. ECG recordings further support adaptive refinement of artifact removal, yielding clean EEG data for subsequent analysis.
Video Stimuli. The 24 fps, 720p video stimuli consist of 20 episodes, each lasting 18 minutes. For video decoding, we downsample the videos to a resolution of 480 × 720 and segment them into 4-second clips (33 frames per clip). Each participant thus contributes 5,400 clips in total, including 4,860 for training (from the first 18 episodes) and 540 for testing (from the final two episodes).
Audio Stimuli. Audio from each video episode is also segmented into corresponding 4-second clips, resulting in the same number of samples as the video clips, maintaining consistency for training and evaluation.
Text Descriptions. To support multimodal decoding experiments, textual descriptions are generated separately for each audio and video clip. Video descriptions are generated using Qwen2.5-VL [Qwen2.5-VL] to facilitate contrastive learning. Audio clips are transcribed using Whisper-large-v3 [whisper], providing rich textual inputs aligned with auditory data for enhanced multimodal analyses.
3.3 Extended Downstream Tasks
Although CineBrain is primarily collected to support multimodal brain-signal-based video decoding, it also provides a solid foundation for a range of extended downstream tasks that may inspire broader multimodal research: (1) Auditory decoding: leveraging both fMRI and EEG signals to reconstruct auditory stimuli. (2) EEG-to-fMRI translation: exploring cross-modal mappings from EEG to fMRI representations, which is particularly meaningful given the high cost and difficulty of fMRI acquisition. (3) Stimulus-to-brain modeling: further extending the framework to predict fMRI or EEG responses directly from video or audio stimuli. Together, these tasks demonstrate the extensibility of CineBrain beyond video reconstruction. We envision CineBrain serving as a comprehensive benchmark for advancing research on how the human brain perceives, integrates, and reconstructs dynamic audiovisual experiences.